# Sylvester-Preconditioned Adaptive-Rank Implicit Time Integrators for Advection-Diffusion Equations with Inhomogeneous Coefficients

### v1

@@ C.3 DIRK Generalization to Multi-Rank Advection-Diffusion Coefficients
We now extend the formulation to the multi-rank case, as described in (4). Using a DIRK method for temporal discretization, the $k$-th stage generalized Sylvester equation becomes:
@@
Using approximations in (25), we arrive at the $k$th stage approximated Sylvester equation:
@@
In order to obtain a reduced equation for the $k$th stage matrix of coefficients $S^{(k)}$, we perform a Galerkin projection $U^{\top}\mathbf{R}_{\mathscr{{L}}}V=0$ to obtain :
@@
Using the operators $\tilde{P}_{1}$ and $\tilde{P}_{2}$, we define the preconditioner mapping:
@@
Algorithm 4 solves the variable advection diffusion equation using DIRK methods.
@@
Tolerances $\epsilon_{\kappa}$, $\epsilon$, ${\epsilon}_{\text{tol}}$;
Maximum iterations max_iter;
@@
2
3for $m=1$ to max_iter do
@@
5        Truncate to tolerance $\epsilon_{\kappa}$;
6        $U_{1}\leftarrow\mathscr{T}_{\epsilon_{\kappa}}\left(\kappa_{m}\left(P_{1},P_{1}^{-1},A^{-1}_{1,1:\ell_{x}},A^{-1}_{3,1:k_{x}},\Phi^{1,y}_{1:\ell_{y}},\Sigma^{1,y}_{1:k_{y}},U_{0}\right)\right)$;
7        $V_{1}\leftarrow\mathscr{T}_{\epsilon_{\kappa}}\left(\kappa_{m}\left(P_{2},P_{2}^{-1},A^{-1}_{2,1:\ell_{y}},A^{-1}_{4,1:k_{y}},\Phi^{2,x}_{1:\ell_{x}},\Sigma^{2,x}_{1:k_{x}},V_{0}\right)\right)$;
8        Compute $\tilde{B}^{(1)}=(U_{1}^{\top}U_{0})S_{0}(V_{1}^{\top}V_{0})^{\top}$;
@@
11               Compute $\tilde{B}^{(k)}=\tilde{B}^{(1)}+\Delta t\sum_{\ell=1}^{k-1}\frac{a_{k\ell}}{a_{\ell\ell}}\left(S^{(\ell)}-\tilde{B}^{(\ell)}\right)$;
12               Solve the reduced equation (65) for $S^{(k)}$ using GMRES with ACS preconditioner (66) to tolerance $\epsilon_{\texttt{GMRES}}$;
13               Compute and store $\frac{1}{a_{kk}}\left(S^{(k)}-\tilde{B}^{(k)}\right)$ and proceed to the next stage;
14
15       Truncate $\{U_{1},S_{1},V_{1}\}\leftarrow\mathcal{T}_{\epsilon}(\{U_{1},S^{(s)},V_{1}\})$ to tolerance $\epsilon$;
@@
17        Compute {_$\{\_,R_{U}\}=\texttt{QR}\left([U_{1},T_{1,1:\ell_{x}}U_{1},\Phi_{1:\ell_{y}}^{1,y}U_{1},T_{3,1:k_{x}}U_{1},\Sigma_{1:k_{y}}^{1,y}U_{1}]\right)$ and
18        {_$\{\_,R_{V}\}=\texttt{QR}\left([V_{1},\Phi_{1:\ell_{x}}^{2,x}V_{1},T_{2,1:\ell_{y}}V_{1},\Sigma_{1:k_{x}}^{2,x}V_{1},T_{4,1:k_{y}}V_{1}]\right)$;
@@
20        Compute the residual $\|\mathbf{R}\|=\left\|R_{U}\text{diag}\left(S_{1}-\tilde{B}^{(s)},-a_{ss}\Delta t\left(I_{\mathcal{R}_{\text{ranks}}}\otimes S_{1}\right)\right)R_{V}^{\top}\right\|$;
@@ 27               Exit loop;
Algorithm 4 s-Stages Adaptive-Rank DIRK Integrator for the Generalized Sylvester Equation. 00footnotetext: The MATLAB implementation of the algorithm is publicly available on GitHub, Adaptive-rank-ACS, https://github.com/helkahza1/Adaptive-rank-ACS

### v2

@@ C.3 DIRK Generalization to Multi-Rank Advection-Diffusion Coefficients
We now extend the formulation to the multi-rank case, as described in (7). Using a DIRK method for temporal discretization, the $k$-th stage GSE becomes:
@@
Using approximations in (62), we arrive at the $k$th stage approximated SE:
@@
In order to obtain a reduced equation for the $k$th stage matrix of coefficients ${\bf S}^{(k)}$, we perform a Galerkin projection $U^{\top}\mathbf{R}_{\mathscr{{L}}}V=0$ to obtain :
@@
Using the operators $\tilde{P}_{1}$ and $\tilde{P}_{2}$, we define the ACS-preconditioner mapping:
@@
Algorithm 5 solves the time-dependent variable advection-diffusion equation using DIRK methods.
@@
Tolerances $\epsilon_{\kappa}$, $\epsilon_{\texttt{GMRES}}$, $\epsilon$, ${\epsilon}_{\text{tol}}$;
@@
2 Compute $l=2+\ell_{x}+\ell_{y}+k_{x}+k_{y}$;
3 Set $U^{(i)}=U_{0}$ and $V^{(i)}=V_{0}$ for $i=1,\dots,l$;
4 Set $U_{1}=U_{0}\quad V_{1}=V_{0}$;
5 while not converged do
@@ 4 Set $U_{1}=U_{0}\quad V_{1}=V_{0}$;
7        Orthogonalize and truncate to tolerance $\epsilon_{\kappa}$;
8        $U_{1},\{U^{(1)}\cdots U^{(l)}\}\leftarrow\kappa_{m}^{\texttt{trunc}}\left(\{P_{1},P_{1}^{-1},A^{-1}_{1,1:\ell_{x}},A^{-1}_{3,1:k_{x}},\Phi^{1,y}_{1:\ell_{y}},\Sigma^{1,y}_{1:k_{y}}\};\{U^{(1)}\cdots U^{(l)}\};U_{1};\epsilon_{\kappa}\right)$;
9        $V_{1},\{V^{(1)}\cdots V^{(l)}\}\leftarrow\kappa_{m}^{\texttt{trunc}}\left(\{P_{2},P_{2}^{-1},A^{-1}_{2,1:\ell_{y}},A^{-1}_{4,1:k_{y}},\Phi^{2,x}_{1:\ell_{x}},\Sigma^{2,x}_{1:k_{x}}\};\{V^{(1)}\cdots V^{(l)}\};V_{1};\epsilon_{\kappa}\right)$;
@@ 7        Orthogonalize and truncate to tolerance $\epsilon_{\kappa}$;
12               Compute $\tilde{{\bf B}}^{(k)}=\tilde{{\bf B}}^{(1)}+\Delta t\sum_{\ell=1}^{k-1}\frac{a_{k\ell}}{a_{\ell\ell}}\left({\bf S}^{(\ell)}-\tilde{{\bf B}}^{(\ell)}\right)$;
13               Solve the reduced equation (59) for ${\bf S}^{(k)}$ using GMRES with ACS preconditioner (60) to tolerance $\epsilon_{\texttt{GMRES}}$;
14               Compute and store $\frac{1}{a_{kk}}\left({\bf S}^{(k)}-\tilde{{\bf B}}^{(k)}\right)$ and proceed to the next stage;
15       Truncate $\{U_{1,\epsilon},S_{1,\epsilon},V_{1,\epsilon}\}\leftarrow\mathcal{T}_{\epsilon}(\{U_{1},S^{(s)},V_{1}\})$ to tolerance $\epsilon$;
@@
17        Compute {_$\{\_,R_{U}\}=\texttt{QR}\left([U_{1},U_{1,\epsilon},T_{1,1:\ell_{x}}U_{1,\epsilon},\Phi_{1:\ell_{y}}^{1,y}U_{1,\epsilon},T_{3,1:k_{x}}U_{1,\epsilon},\Sigma_{1:k_{y}}^{1,y}U_{1,\epsilon}]\right)$ and
18        {_$\{\_,R_{V}\}=\texttt{QR}\left([V_{1},V_{1,\epsilon},\Phi_{1:\ell_{x}}^{2,x}V_{1,\epsilon},T_{2,1:\ell_{y}}V_{1,\epsilon},\Sigma_{1:k_{x}}^{2,x}V_{1,\epsilon},T_{4,1:k_{y}}V_{1,\epsilon}]\right)$ ;
@@
20        Compute the residual $\|\mathbf{R}\|=\left\|R_{U}\text{diag}\left(-\tilde{B}^{(s)},S_{1,\epsilon},-\Delta t\left(I_{\mathcal{R}_{\text{ranks}}}\otimes S_{1,\epsilon}\right)\right)R_{V}^{\top}\right\|$;
@@ 28               Exit loop;
Algorithm 5 s-Stages Adaptive-Rank DIRK Integrator for the GSE.
Appendix D Three-Dimensional Formulation
We outline our approach to 3D next. In 3D, we replace the SVD algorithm with a tensor Tucker decomposition, specifically tailored for the three (and higher) dimension tensor-product grids.
D.1 Tucker Tensor Equations
We discretize the 3D time-dependent advection-diffusion equation on a three-dimensional tensor-product spatial grid with $N_{1}$, $N_{2}$, and $N_{3}$ grid points in the $x$, $y$, and $z$ directions, respectively. Here and in the following, we let ${\bf F}\in\mathbb{R}^{N_{1}\times N_{2}\times N_{3}}$ denote a three-dimensional tensor. The mode-$n$ product of ${\bf F}$ with a matrix $A_{n}\in\mathbb{R}^{I\times N_{n}}$, for $n=1,2,3$, is defined as:
${\bf F}\times_{1}A_{1}\in\mathbb{R}^{I\times N_{2}\times N_{3}},\,{\bf F}\times_{2}A_{2}\in\mathbb{R}^{N_{1}\times I\times N_{3}},\,{\bf F}\times_{3}A_{3}\in\mathbb{R}^{N_{1}\times N_{2}\times I}.$
The resulting tensor in each case has its $n$-th dimension replaced by the number of rows of $A_{n}$. The entries of the mode-$n$ product are computed as:
$({\bf F}\times_{n}A_{n})_{i_{1},\ldots,i_{n-1},j,i_{n+1},\ldots,i_{3}}=\sum_{i_{n}=1}^{N_{n}}{\bf F}_{i_{1},\ldots,i_{n},\ldots,i_{3}}\cdot(A_{n})_{j,i_{n}}.$
For a more thorough review of tensor operations, we refer the reader to [22]. The discretization of the 3D advection-diffusion equation using finite-differences in space combined with the backward Euler method in time, yields the following GSE:
${\bf R}_{\mathscr{L}}=\mathbf{F}_{1}-\Delta t\mathscr{L}({\bf F}_{1})-\mathbf{F}_{0}=\mathbf{0},$ (61)
where ${\bf F}_{0}$ denotes the initial tensor condition at time $t_{0}$, and ${\bf F}_{1}$ is the sought tensor solution at time $t_{1}$, and,
$\mathscr{L}:\mathbf{F}\mapsto$ $\sum_{i=1}^{\ell_{x}}\mathbf{F}\times_{1}T_{1,i}\times_{2}\Phi_{i}^{2,x}\times_{3}\Phi_{i}^{3,x}+\sum_{i=1}^{\ell_{y}}\mathbf{F}\times_{1}\Phi_{i}^{1,y}\times_{2}T_{2,i}\times_{3}\Phi_{i}^{3,y}+\sum_{i=1}^{\ell_{z}}\mathbf{F}\times_{1}\Phi_{i}^{1,z}\times_{2}\Phi_{i}^{2,z}\times_{3}T_{3,i}$
$+$ $\sum_{i=1}^{k_{x}}\mathbf{F}\times_{1}T_{4,i}\times_{2}\Sigma_{i}^{2,x}\times_{3}\Sigma_{i}^{3,x}+\sum_{i=1}^{k_{y}}\mathbf{F}\times_{1}\Sigma_{i}^{1,y}\times_{2}T_{5,i}\times_{3}\Sigma_{i}^{3,y}+\sum_{i=1}^{k_{z}}\mathbf{F}\times_{1}\Sigma_{i}^{1,z}\times_{2}\Sigma_{i}^{2,z}\times_{3}T_{6,i}.$
The operators $T$ correspond to the composition of advection and diffusion difference operators with diagonal matrices, analogous to those defined in Section 2. Similarly, the operators $\Phi$ and $\Sigma$ are analogous to the diagonal matrices for diffusion and advection coefficients, respectively, used for pointwise multiplication, as defined in Section 2. To obtain an averaged (approximate) mapping for preconditioning purposes, we define the average coefficients:
$\alpha_{i}^{2,x}$ $=\text{avg}(\Phi_{i}^{2,x}),\,\alpha_{i}^{3,x}$ $=\text{avg}(\Phi_{i}^{3,x}),\,\alpha_{i}^{1,y}$ $=\text{avg}(\Phi_{i}^{1,y}),\,\alpha_{i}^{3,y}$ $=\text{avg}(\Phi_{i}^{3,y}),\,\alpha_{i}^{1,z}$ $=\text{avg}(\Phi_{i}^{1,z}),\,\alpha_{i}^{2,z}$ $=\text{avg}(\Phi_{i}^{2,z}),$ (62)
$\gamma_{i}^{2,x}$ $=\text{avg}(\Sigma_{i}^{2,x}),\,\gamma_{i}^{3,x}$ $=\text{avg}(\Sigma_{i}^{3,x}),\,\gamma_{i}^{1,y}$ $=\text{avg}(\Sigma_{i}^{1,y}),\,\gamma_{i}^{3,y}$ $=\text{avg}(\Sigma_{i}^{3,y}),\,\gamma_{i}^{1,z}$ $=\text{avg}(\Sigma_{i}^{1,z}),\,\gamma_{i}^{2,z}$ $=\text{avg}(\Sigma_{i}^{2,z}).$
which yields the three-dimensional approximate operator $\tilde{\mathscr{L}}$ as:
[…]
$+\mathbf{F}\times_{2}\bigg{(}\sum_{j=1}^{\ell_{y}}\alpha_{j}^{1,y}\alpha_{j}^{3,y}T_{2,j}+\sum_{l=1}^{k_{y}}\gamma_{l}^{1,y}\gamma_{l}^{3,y}T_{5,l}\bigg{)}$
$+\mathbf{F}\times_{3}\bigg{(}\sum_{k=1}^{\ell_{z}}\alpha_{k}^{1,z}\alpha_{k}^{2,z}T_{3,k}+\sum_{l=1}^{k_{z}}\gamma_{l}^{1,z}\gamma_{l}^{2,z}T_{6,l}\bigg{)}.$ (63)
Using this approximate operator, and letting ${\mathcal{R}_{\text{ranks}}}=r_{x}+r_{y}+r_{z}+\ell_{x}+\ell_{y}+\ell_{z}$, we then define the averaged three-dimensional approximate SE of the form:
${\bf F}_{1}\times_{1}P_{1}+{\bf F}_{1}\times_{2}P_{2}+{\bf F}_{1}\times_{3}P_{3}={\bf F}_{0},$ (64)
with
$\small P_{1}$ $=\sum_{i=1}^{\ell_{x}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-\Delta t\alpha_{i}^{2,x}\alpha_{i}^{3,x}T_{1,i}\right)}_{{A_{1,i}}}+\sum_{k=1}^{k_{x}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-\Delta t\gamma_{k}^{2,x}\gamma_{k}^{3,x}T_{4,k}\right)}_{{A_{4,k}}}=\sum_{i=1}^{\ell_{x}}{A_{1,i}}+\sum_{k=1}^{k_{x}}{A_{4,k}},$
$P_{2}$ $=\sum_{j=1}^{\ell_{y}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\alpha_{j}^{1,y}\alpha_{j}^{3,y}T_{2,j}\right)}_{A_{2,i}}+\sum_{l=1}^{k_{y}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\gamma_{l}^{1,y}\gamma_{l}^{3,y}T_{5,l}\right)}_{A_{5,i}}=\sum_{j=1}^{\ell_{y}}{A_{2,j}}+\sum_{l=1}^{k_{y}}{A_{5,l}},$
$P_{3}$ $=\sum_{j=1}^{\ell_{z}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{3}}-\Delta t\alpha_{k}^{1,z}\alpha_{k}^{2,z}T_{3,k}\right)}_{A_{3,i}}+\sum_{l=1}^{k_{z}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\gamma_{l}^{1,z}\gamma_{l}^{2,z}T_{6,l}\right)}_{A_{6,i}}=\sum_{j=1}^{\ell_{z}}{A_{3,j}}+\sum_{l=1}^{k_{z}}{A_{6,l}}.$
These one-dimensional matrix operators, analogous to those in (26), are used (i) to construct xKrylov subspaces for the Tucker basis factors and (ii) to precondition the solution of the three-dimensional GSE arising from the Galerkin condition on the residual in Eq. (61).
D.2 GMRES-ACS Solution of the Reduced GSE in 3D
We assume the existence of a low-rank Tucker factorization of the initial condition, $\mathbf{F}_{0}\in\mathbb{R}^{N_{1}\times N_{2}\times N_{3}}$, which evolves to an updated solution $\mathbf{F}_{1}\in\mathbb{R}^{N_{1}\times N_{2}\times N_{3}}$ at time $t^{(1)}=t^{(0)}+\Delta t$, also in the low-rank Tucker format. This is represented as:
[…]
where $U_{0}\in\mathbb{R}^{N_{1}\times r^{x}_{0}}$, $V_{0}\in\mathbb{R}^{N_{2}\times r^{y}_{0}}$, $W_{0}\in\mathbb{R}^{N_{3}\times r^{z}_{0}}$, $U_{1}\in\mathbb{R}^{N_{1}\times r^{x}_{1}}$, $V_{1}\in\mathbb{R}^{N_{2}\times r^{y}_{1}}$, and $W_{1}\in\mathbb{R}^{N_{3}\times r^{z}_{1}}$ are orthonormal bases for their respective dimensions, and $S_{0}\in\mathbb{R}^{r^{x}_{0}\times r^{y}_{0}\times r^{z}_{0}}$ and $S_{1}\in\mathbb{R}^{r^{x}_{1}\times r^{y}_{1}\times r^{z}_{1}}$ are core tensors at the corresponding times. The bases $U_{1}$, $V_{1}$, and $W_{1}$ are constructed from orthonormalization of dimension-wise xKrylov, as discussed in 3.2
Using a Galerkin projection on the residual, $\mathbf{R}_{\mathscr{L}}\times_{1}U_{1}^{\top}\times_{2}V_{1}^{\top}\times_{3}W_{1}^{\top}$, we obtain a reduced-size GSE for the core tensor ${\bf S}_{1}$ as follows:
$\mathbf{S}_{1}-\Delta t\mathscr{L}(\mathbf{S}_{1})\times_{1}U_{1}^{\top}\times_{2}V_{1}^{\top}\times_{3}W_{1}^{\top}=\mathbf{S}_{0}\times_{1}U_{1}^{\top}U_{0}\times_{2}V_{1}^{\top}V_{0}\times_{3}W_{1}^{\top}W_{0}.$ (66)
This projected system has dimensions $r_{x}\times r_{y}\times r_{z}$, where $r_{x}$, $r_{y}$, and $r_{z}$ denote the dimensions of the truncated xKrylov subspaces in the $x$, $y$, and $z$ directions, respectively, and will be solved using the GMRES-ACS preconditioner.
Eq. (66) is solved iteratively using GMRES, with $\mathscr{L}$-operator application cast as a series of tensor-matrix multiplications, along with the ACS preconditioning operator given by:
$\mathcal{P}:\mathbf{S}\mapsto\mathbf{S}\times_{1}\tilde{P}_{1}+\mathbf{S}\times_{2}\tilde{P}_{2}+\mathbf{S}\times_{3}\tilde{P}_{3},$ (67)
where $\tilde{P}_{i}=P_{i}\times_{1}U_{1}^{\top}\times_{2}V_{1}^{\top}\times_{3}W_{1}^{\top}$ denotes the projection of operators in Eq. (64), analogous to those introduced in Section 3. This operator is used to precondition each GMRES iteration of the inner system solve. The solution of the projected SE has a computational cost of $\mathcal{O}(r^{4})$ [9, 41]. The tensor-times-matrix operations, executed using the ttm command in the Tensor Toolbox, incur a cost of $\mathcal{O}(r^{4})$. The total complexity in three dimensions is therefore $\mathcal{O}(Nr^{2}+r^{4})$.